OpenStudy (anonymous):

a small manufacturing and whole sale company estimates that its daily profit P in doars is P(x) = 10x - 0.01x raised to the power of 2. a) How many items must be sold daily for the maximum profit? b) What is the maximum daiy profit?

OpenStudy (anonymous):

P(x) = 10x -0.01x^2 P'(x)= 10 -0.02x =0 for min/max 0.02x= 10 x= 500 then sub this into the formula P(500) = 10(500) - (0.01)(500^2) = $2500

OpenStudy (anonymous):

thanks